OpenStudy (anonymous):

Write the inverse variation equation of the following. Y varies inversely with x and y = 9 when x = 14. Leave no spaces in your equation.

OpenStudy (anonymous):

\( Y = \frac K x \)

OpenStudy (anonymous):

its y=1/x

OpenStudy (anonymous):

Im confused :( how do i write it as a ordered pair and wheres the numbers?

OpenStudy (anonymous):

u can write as xy =1

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i mean its bcam as a equation in xy =1

OpenStudy (anonymous):

Sorry I accidentally deleted my post. \[\Large \begin{array}{l} y = \frac{k}{x}\\ 9 = \frac{k}{{14}}\\ k = ? \end{array}\] Calculate what k is.

OpenStudy (anonymous):

bt how can we identify .....in this case there is three variable i.e x,y & k also ...

OpenStudy (anonymous):

i think it must be only in inversely i.e in form of 1/x only right ?

OpenStudy (anonymous):

It's inversely proportional. Meaning, yes, as x increases y decreases and vice versa, which will happen if y = 1/x. But when x = 14, y = 9, meaning the constant numerator is not 1; otherwise y = 1/14.